OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different? PLEASE PLEASE PLEASE HELP

OpenStudy (anonymous):

part1: x2=16 x=+-4 x2+2x-24=(x+6)(x-4) x=-6, x=4 part2: solutions are the point where the graph crosses x axis

OpenStudy (anonymous):

THANK YOU!!! part 3?

OpenStudy (precal):